OpenStudy (theflash12345):

How do I solve |4-2y|+5=9

Vocaloid (vocaloid):

start by subtracting 5 from both sides |4 - 2y| = 4 then we write out both cases 1. 4 - 2y = 4 2. 4 - 2y = -4 then solve #1 and #2 for y. you will get 2 different y values

OpenStudy (theflash12345):

Ok then what? Or is this it

Vocaloid (vocaloid):

solve the two equations i gave you

Vocaloid (vocaloid):

4 - 2y = 4 y = ?

OpenStudy (theflash12345):

I did

Vocaloid (vocaloid):

that's it

OpenStudy (theflash12345):

Ok thanks